what is the slope [and the y-intercept] of y-5=1/3(x-6)
Answer by TimothyLamb(4379)

First, get the equation into slope-intercept form: y = mx + b
---
by algebra like so:
---
y - 5 = (1/3)x - (6/3)
y - 5 = (1/3)x - 2
y = (1/3)x - 2 + 5
y = (1/3)x + 3
---
Answer:
Slope (m) = 1/3
Y-intercept (b) = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php